With a population of 15 and a sample size of 3, what is the total number of samples that are possible using:


Theoretical Sampling?




Experimental Sampling?




1
Expert's answer
2021-02-24T06:04:13-0500

Here N=15 and n=3

(a) Theoretical sampling: Theoretical sampling is process of data collection for generating theory.

Number of possible sample"\Rightarrow N^n =" 153=3375


(b)Experimental sampling: In experimental sampling using simple random sampling.

Suppose sample of size n is drawn from population of size N, then there will be "^{N}C_n" possible samples.

"\therefore" Possible number of samples = "^{N}C_n= ^{15}C_3 = \dfrac{15!}{3!\times12!}"

"\Rightarrow \dfrac{15\times14\times13}{(3\times2\times1)}=" 455